# Static Word Embeddings for Sentence Semantic Representation

### v1

@@ Abstract
We propose new static word embeddings optimised for sentence semantic representation. We first extract word embeddings from a pre-trained Sentence Transformer, and improve them with sentence-level principal component analysis, followed by either knowledge distillation or contrastive learning. During inference, we represent sentences by simply averaging word embeddings, which requires little computational cost. We evaluate models on both monolingual and cross-lingual tasks and show that our model substantially outperforms existing static models on sentence semantic tasks, and even rivals a basic Sentence Transformer model (SimCSE) on some data sets. Lastly, we perform a variety of analyses and show that our method successfully removes word embedding components that are irrelevant to sentence semantics, and adjusts the vector norms based on the influence of words on sentence semantics.
@@ 1 Introduction
In natural language processing (NLP), it has been a major topic to learn fixed-length embeddings that represent sentence semantics. To achieve this goal, most existing methods fine-tune pre-trained language models (LMs) such as BERT Devlin et al. (2019) using labelled data of various NLP tasks (e.g. natural language inference). Recent work further employs large language models (LLMs) such as LLaMA 3 AI@Meta (2024) and shows that scaling up models can yield better performance BehnamGhader et al. (2024). However, this approach requires a large amount of computational cost, making it difficult to process billions of sentences cost-efficiently or deploy models on resource-constrained devices such as smartphones. As such, it is crucial to explore more cost-efficient models that work well without GPUs, and yet this direction remains largely overlooked in the research community.
In this work, we propose new static word embeddings (SWEs) — both monolingual and cross-lingual ones — that can represent sentence semantics fairly well with little computational cost. Specifically, our method first extracts SWEs from a pre-trained Sentence Transformer Vaswani et al. (2017); Reimers and Gurevych (2019), and improves them with sentence-level principal component analysis (PCA) followed by either knowledge distillation or contrastive learning. During inference, we encode sentences by simply looking up SWEs and taking their average, which can be done comfortably with limited resources. Our main contributions are: (1) we propose new SWEs that substantially outperform existing SWEs on a semantic textual similarity (STS) task, and even surpass SimCSE Gao et al. (2021) on Massive Text Embedding Benchmark Muennighoff et al. (2023) in English; (2) we propose cross-lingual SWEs optimised for sentence representation, which achieve surprisingly good performance on translation retrieval even between distant languages (e.g. 94.6% in F1 scores on English–Chinese); and (3) we conduct a variety of analyses and provide new insights into how to obtain SWEs suitable for sentence representation. Our SWEs will be available at https://github.com/twadada/swe4semantics.
@@ 2 Related Work
Most recent text embedding models are built on pre-trained Transformer LMs. For instance, GTE Li et al. (2023) fine-tunes BERT with contrastive learning on labelled data from various sources (e.g. STS data sets). Later, it is extended to the multilingual model called mGTE Zhang et al. (2024), which is trained on both monolingual and cross-lingual data sets. LLM2Vec BehnamGhader et al. (2024) scales up the backbone model and fine-tunes LLMs such as LLaMA with contrastive learning. Similarly, NV-Embed Lee et al. (2025) fine-tunes Mistral-7B Jiang et al. (2023) with contrastive learning and also trains an additional attention layer on top. While these models achieve impressive results, they are computationally expensive and inefficient for real-world applications. To mitigate this, several methods to train compact models are proposed such as knowledge distillation Hinton et al. (2015); Xu et al. (2023), but they are mostly applied to Transformer models, which are designed with GPU acceleration in mind.
@@
On the other hand, little attention has been paid to building cost-efficient models (esp. cross-lingual ones) that can be easily run on CPUs. Although such models inevitably lag behind heavy-weight models in accuracy, they offer significant advantages in cost and runtime efficiency, and can be applied to, for instance, quickly mining semantically similar content from billions of posts on social media. One traditional model in this direction is Sent2Vec Pagliardini et al. (2018), which extends the CBOW algorithm Mikolov et al. (2013a) and represents sentences by averaging static word and bigram embeddings. Very recently, two static word embedding models called WordLlama Miller (2024) and Model2Vec Tulkens and van Dongen (2024) have been released from different GitHub projects (MIT license), [^1] https://github.com/dleemiller/WordLlama and https://github.com/MinishLab/model2vec, respectively. and demonstrate strong performance on sentence-level tasks in English. Crucially, however, they do not publish a paper or detailed technical report on these models, leaving their tuning processes unclear and lacking the analyses of their embeddings (unlike our work). Based on their code and documentations, both models extract SWEs from Transformer models. Specifically, WordLlama extracts SWEs from the input layer of LLMs like LLaMa 3 AI@Meta (2024), and refines them with contrastive learning on various labelled data such as STS and text retrieval data sets. Model2Vec extracts SWEs from a Transformer text encoder called BGE-base Xiao et al. (2023), by feeding each word $w$ in a vocabulary into BGE without context (i.e. “[CLS], $w$, [SEP]”) and using the output as the embedding of $w$. It then fine-tunes the SWEs by minimising the mean cosine distance between the sentence embeddings produced by BGE and the average of the SWEs for the same input text. To reduce the dimensionality, both WordLlama and Model2Vec perform word-level PCA as a post-processing step, and Model2Vec also applies the smooth inverse frequency (SIF) heuristic Arora et al. (2017), which multiplies the norm of $w$ by $\tfrac{\alpha}{\alpha+\mathrm{p}(w)}$, where $\mathrm{p}(w)$ denotes the uni-gram probability of $w$ and $\alpha=0.001$.
@@ 3 Methodology
Our proposed method obtains SWEs in three steps: (1) embedding extraction from Sentence Transformer (Sec. 3.1); (2) dimensionality reduction with sentence-level PCA (Sec. 3.2); and (3) embedding refinement with knowledge distillation or contrastive learning (Sec. 3.3). Only the last step involves the training (fine-tuning) of SWEs.
@@ 3.1 Embedding Extraction
We first extract SWEs from a pre-trained Sentence Transformer (ST). Specifically, to obtain the embedding of a word $w\in V$ (denoted as $\mathrm{E}(w)$), we sample $N=100$ sentences $z^{(w)}_{i}~{}(i=1,2,...,N)$ that contain $w$ from a large corpus. Then, we encode each sentence $z^{(w)}_{i}$ using ST, which produces $|z^{(w)}_{i}|$ vectors at the last Transformer layer ($|z^{(w)}_{i}|$ denotes the number of tokens in $z^{(w)}_{i}$). We then retrieve the vector at the position of $w$ (denoted as $\mathrm{ST}(w,z^{(w)}_{i})$), and obtain $\mathrm{E}(w)$ as:
@@
This way, we decontextualise the embeddings and obtain the SWE of $w$. Furthermore, we can also generate cross-lingual SWEs using multilingual ST, as it embeds different languages in the same space. For the vocabulary $V$, we include more words than the ST’s original vocabulary, and if $w$ is tokenised into subwords, we average its subword embeddings to obtain $\mathrm{ST}(w,z^{(w)}_{i})$ in Eqn. (1).
The idea of extracting SWEs from Transformers has been explored in previous studies Ethayarajh (2019); Bommasani et al. (2020); Wada et al. (2022), which show that SWEs extracted from masked LMs perform well on word-level tasks (e.g. lexical substitution). However, we find that these embeddings are not effective for sentence-level tasks, and hence we refine them in the following steps.
@@ 3.2 Sentence-level PCA
Specifically, we randomly sample $M=100\mathrm{k}$ sentences from those used in Section 3.1, and represent each sentence by averaging SWEs. When training cross-lingual SWEs, we sample $M$ sentences from each language and concatenate them. Given the sentence embedding matrix $X\in\mathbb{R}^{M\times d}$ (or $\in\mathbb{R}^{2M\times d}$) and its mean vector $\bar{X}\in\mathbb{R}^{d}$, we apply PCA to the centred matrix $X-\bar{X}$ and obtain the transformation matrix $W$. This matrix transforms the embedding space onto new axes that effectively capture the variance of the sentence representations. Importantly, since we represent each sentence by averaging SWEs, we can pre-transform the SWE $\mathrm{E}(w)$ into $\mathrm{\hat{E}}(w)=W^{\top}(\mathrm{E}(w)-\bar{X})$ for each word $w\in V$ once and save $\mathrm{\hat{E}}(w)$ in memory, as $W^{\top}(\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{E}(w^{\prime})-\bar{X})=\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{\hat{E}}(w^{\prime})$.
To reduce the dimensionality using PCA, a common approach is to keep the principal components (PCs) with the $d^{\prime}(<d)$ largest eigenvalues. However, we instead discard the first $r$ PCs, and keep the $(r+1)$-th to $(r+d^{\prime})$-th components. This is based on the previous finding that removing a few dominant PCs from SWEs can counter-intuitively improve the embedding quality Mu and Viswanath (2018), and we actually get positive results in our experiments especially in cross-lingual tasks (e.g. +49.4% in F1 scores). Intriguingly, our analysis (Sec. 5.2) reveals that it helps remove information not relevant to sentence semantics. Following Mu and Viswanath (2018), we set $r$ to $\lfloor\frac{d}{100}\rfloor$, and refer to this method as ABTT (All-But-The-Top).
@@ 3.3 Embedding Refinement
Lastly, we refine the dimensionality-reduced embeddings after PCA. For monolingual SWEs, we perform teacher-student knowledge distillation, where we fine-tune our SWEs (= student) to reproduce the sentence similarities calculated by ST (= teacher), which is freezed during distillation; hence, it is not computationally expensive and can be done without labelled data (unlike WordLlama). Formally, for a mini-batch of $K=128$ sentences, we minimise the following loss $\mathcal{L}$:
@@
where $\mathbf{S},\mathbf{T}\in\mathbb{R}^{K\times K}$ denote the similarity matrices calculated by the teacher and student models, with the value at the $i$-th row and $j$-th column corresponding to the cosine similarity between the $i$-th and $j$-th sentences in the mini-batch. The temperature $\tau$ is set to $0.05$ as commonly done in contrastive learning Gao et al. (2021). Note that the student model represents each sentence by averaging SWEs obtained in Section 3.2. Since the teacher and student models calculate $S$ and $T$ independently, their embedding spaces can be different and that makes it possible to fine-tune the dimensionality-reduced embeddings after PCA. In contrast, Model2Vec minimises the cosine distance of embeddings output by ST and SWEs, requiring both models to share the same embedding space; hence, it applies (word-level) PCA after fine-tuning, which can diminish the benefits of fine-tuning and hence relies on the SIF heuristic to improve performance. Section 5.3 shows that our method yields a similar effect to SIF, but in a more refined manner.
For cross-lingual SWEs, we find that performing knowledge distillation for each language leads to worse results on cross-lingual tasks, as it does not encourage SWEs to be mapped into the same cross-lingual space. Therefore, we instead make use of a parallel corpus as a source of knowledge, and fine-tune the cross-lingual SWEs in languages $\ell 1$ and $\ell 2$ by minimising the following loss $\mathcal{L}_{CL}$:
@@ 4.1 Our Models
We train our monolingual (English) SWEs using GTE-base Li et al. (2023), which achieves strong performance on Massive Text Embedding Benchmark (MTEB) among BERT-sized models. We sample sentences from a large-scale unlabelled corpus (CC-100 Wenzek et al. (2020); Conneau et al. (2020)) and use them in each step of our method. For the vocabulary $V$, we include the 150k most frequent words (case-insensitive) in CC-100. To generate cross-lingual SWEs, we use mGTE Zhang et al. (2024), the multilingual version of GTE trained on 75 languages. We train cross-lingual embeddings for three language pairs, namely English–German (en–de), English–Chinese (en–zh), and English–Japanese (en–ja). We sample short sentences from large-scale parallel corpora (CC-Matrix Schwenk et al. (2021)) and use them to extract SWEs in Section 3.1 and calculate the loss in Eqn. (4). For en–ja, we also use JParaCrawl v3.0 Morishita et al. (2020, 2022) to augment data. For $V$ in English and German, we choose the 150k most frequent words (case-sensitive) in the corresponding corpora, and 30k most frequent tokens (which can be a subword or consist of multiple words) in Japanese and Chinese. [^3] This is because mGTE’s tokeniser segments sentences into tokens without performing word segmentation in both languages (which have no whitespace word boundary).
In Section 3.2, we reduce the dimensionality to $d^{\prime}\in\{256,512\}$, [^4] In most cases, 99% of the variance can be explained with $d^{\prime}=512$ when the original dimensionality $d$ is $768$. and in Section 3.3 we fine-tune SWEs using Adam Kingma and Ba (2015) with the learning rate of 0.001 for 30k steps, and apply early stopping based on the validation loss; the train and validation data consist of sentences sampled from those used in Section 3.1. During inference, we tokenise the text $z$ into words (or tokens in Japanese and Chinese) and strip punctuation, and average SWEs to represent $z$. When a word is not in $V$, we tokenise it into subwords using (m)GTE’s tokeniser, and truncate it by removing the last subword until the remaining string is found in $V$ [^5] For instance, if tokeniser is out of vocabulary, we split it into subwords (e.g. token #ise #r), and strip the last subword(s). Then, if tokenise (or token) is in $V$, we use its embedding.

### v2

@@ Abstract
We propose new static word embeddings optimised for sentence semantic representation. We first extract word embeddings from a pre-trained Sentence Transformer, and improve them with sentence-level principal component analysis, followed by either knowledge distillation or contrastive learning. During inference, we represent sentences by simply averaging word embeddings, which requires little computational cost. We evaluate models on both monolingual and cross-lingual tasks and show that our model substantially outperforms existing static models on sentence semantic tasks, and even surpasses a basic Sentence Transformer model (SimCSE) on a text embedding benchmark. Lastly, we perform a variety of analyses and show that our method successfully removes word embedding components that are not highly relevant to sentence semantics, and adjusts the vector norms based on the influence of words on sentence semantics.
@@ 1 Introduction
In natural language processing (NLP), it has been a major topic to learn fixed-length embeddings that represent sentence semantics. To achieve this goal, most existing methods fine-tune pre-trained language models (LMs) such as BERT Devlin et al. (2019) using labelled data of various NLP tasks (e.g. natural language inference). Recent work further employs large language models (LLMs) such as LLaMA 3 AI@Meta (2024) as backbone models and shows that scaling up models can yield better performance BehnamGhader et al. (2024); Lee et al. (2025). However, this approach requires a large amount of computational cost, making it difficult to process billions of sentences cost-efficiently or deploy models on resource-constrained devices such as smartphones. As such, it is crucial to explore more cost-efficient models that work well without GPUs, and yet this direction remains largely overlooked in the research community.
In this work, we propose new static word embeddings (SWEs) — both monolingual and cross-lingual ones — that can represent sentence semantics fairly well with little computational cost. Specifically, our method first extracts SWEs from a pre-trained Sentence Transformer Vaswani et al. (2017); Reimers and Gurevych (2019), and improves them with sentence-level principal component analysis (PCA), followed by either knowledge distillation or contrastive learning. During inference, we encode sentences by simply looking up SWEs and taking their average, which can be done comfortably with limited resources. Our main contributions are: (1) we propose new SWEs that substantially outperform existing SWEs on a semantic textual similarity (STS) task, and even surpass SimCSE Gao et al. (2021) on Massive Text Embedding Benchmark Muennighoff et al. (2023) in English; (2) we propose cross-lingual SWEs optimised for sentence representation, which achieve surprisingly good performance on translation retrieval even between distant languages (e.g. 94.6% in F1 scores on English–Chinese); and (3) we conduct a variety of analyses and provide new insights into how to obtain SWEs suitable for sentence representation. We will release our code and models at https://github.com/twadada/swe4semantics.
@@ 2 Related Work
Most recent text embedding models are built on pre-trained Transformer LMs. For instance, GTE Li et al. (2023) fine-tunes BERT with contrastive learning on labelled data retrieved from various sources (e.g. STS data sets). Later, it is extended to the multilingual model called mGTE Zhang et al. (2024), which is trained on both monolingual and cross-lingual data sets. LLM2Vec BehnamGhader et al. (2024) scales up the backbone model and fine-tunes LLMs such as LLaMA with contrastive learning. Similarly, NV-Embed Lee et al. (2025) fine-tunes Mistral-7B Jiang et al. (2023) with contrastive learning and also trains an additional attention layer on top. While these models achieve impressive results, they are computationally expensive and inefficient for real-world applications. To mitigate this, several methods to train compact models are proposed (e.g. knowledge distillation Hinton et al. (2015); Xu et al. (2023)), but they are mostly applied to Transformer models, which are designed with GPU acceleration in mind.
@@
On the other hand, little attention has been paid to building cost-efficient models (esp. cross-lingual ones) that can be easily run on CPUs. Although such models inevitably lag behind heavy-weight models in accuracy, they offer significant advantages in cost and runtime efficiency, and can be applied to, for instance, quickly mining semantically similar content from billions of posts on social media. One traditional model in this direction is Sent2Vec Pagliardini et al. (2018), which extends the CBOW algorithm Mikolov et al. (2013a) and represents sentences by averaging static word and bigram embeddings. Very recently, although not published in papers, two static word embedding models called WordLlama Miller (2024) and Model2Vec Tulkens and van Dongen (2024) have been released from different GitHub projects (MIT license), [^1] https://github.com/dleemiller/WordLlama and https://github.com/MinishLab/model2vec, respectively. and demonstrate strong performance on sentence-level tasks in English. Based on their code and documentations, both models extract SWEs from Transformer models. Specifically, WordLlama extracts SWEs from the input layer of LLMs like LLaMa 3 AI@Meta (2024), and refines them with contrastive learning on various labelled data such as STS and text retrieval data sets. Model2Vec extracts SWEs from a Transformer text encoder called BGE-base Xiao et al. (2023), by feeding each word $w$ in a vocabulary into BGE without context (i.e. “[CLS], $w$, [SEP]”) and using the output as the embedding of $w$. It then fine-tunes the SWEs by minimising the mean cosine distance between the sentence embeddings produced by BGE and the average of the SWEs for the same input text. To reduce the dimensionality, both WordLlama and Model2Vec apply PCA to a word embedding matrix as a post-processing step, and Model2Vec also applies the smooth inverse frequency (SIF) heuristic Arora et al. (2017), which multiplies the norm of $w$ by $\tfrac{\alpha}{\alpha+\mathrm{p}(w)}$, where $\mathrm{p}(w)$ denotes the uni-gram probability of $w$ and $\alpha=0.001$.
@@ 3 Methodology
Our proposed method obtains SWEs in three steps: (1) embedding extraction from Sentence Transformer (Sec. 3.1); (2) dimensionality reduction with sentence-level PCA (Sec. 3.2); and (3) embedding refinement with knowledge distillation or contrastive learning (Sec. 3.3). Note that only the last step involves the training (fine-tuning) of SWEs.
@@ 3.1 Embedding Extraction
[…]
@@
This way, we decontextualise the embeddings and obtain the SWE of $w$. Furthermore, we can also generate cross-lingual SWEs using multilingual ST, as it embeds different languages in the same space. For the vocabulary $V$, we include more words than those in the ST’s original vocabulary, and if $w$ is tokenised into subwords, we average its subword embeddings to obtain $\mathrm{ST}(w,z^{(w)}_{i})$ in Eqn. (1).
The idea of extracting SWEs from Transformers has been explored in previous studies Ethayarajh (2019); Bommasani et al. (2020); Wada et al. (2022), which show that SWEs extracted from masked LMs perform well on word-level tasks (e.g. lexical substitution). However, we find that these embeddings are not effective for sentence-level tasks (as we will show in our ablation studies in Section 5.1), and hence we further refine them in the following steps.
@@ 3.2 Sentence-level PCA
Specifically, we randomly sample $M=100\mathrm{k}$ sentences from those used in Section 3.1, and represent each sentence by averaging SWEs. When training cross-lingual SWEs, we sample $M$ sentences from $L$ languages and concatenate them; in our experiments, we train bilingual SWEs (i.e. $L=2$) unless otherwise specified. Given the sentence embedding matrix $X\in\mathbb{R}^{LM\times d}$ (with $L=1$ denoting monolingual SWEs) and its mean vector $\bar{X}\in\mathbb{R}^{d}$, we apply PCA to the centred matrix $X-\mathbf{1}_{M}\bar{X}^{\top}$, where $\mathbf{1}_{M}\in\mathbb{R}^{M}$ is a vector of ones, and then obtain the transformation matrix $W$. This matrix transforms the embedding space onto new axes that effectively capture the variance of the sentence representations. Importantly, since we represent each sentence by averaging SWEs, we can pre-transform the SWE $\mathrm{E}(w)$ into $\mathrm{\hat{E}}(w)=W^{\top}(\mathrm{E}(w)-\bar{X})$ for each word $w\in V$ once and save $\mathrm{\hat{E}}(w)$ in memory, as $W^{\top}(\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{E}(w^{\prime})-\bar{X})=\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{\hat{E}}(w^{\prime})$.
To reduce the dimensionality using PCA, a common approach is to keep the principal components (PCs) with the $d^{\prime}(<d)$ largest eigenvalues. However, we instead discard the first $r$ PCs, and keep the $(r+1)$-th to $(r+d^{\prime})$-th components. This is based on the previous finding that removing a few dominant PCs from SWEs can counter-intuitively improve the embedding quality Mu and Viswanath (2018), and we actually get positive results in our experiments, especially in cross-lingual tasks (e.g. +49.4% in F1 scores). Intriguingly, our analyses in Section 5.2 reveal that it helps remove information that is not very relevant to sentence semantics. Following Mu and Viswanath (2018), we set $r$ to $\lfloor\frac{d}{100}\rfloor$, and refer to this method as ABTT (All-But-The-Top).
@@ 3.3 Embedding Refinement
Lastly, we fine-tune the dimensionality-reduced embeddings obtained in Section 3.2 to further optimise them for sentence representation. For monolingual SWEs, we perform teacher-student knowledge distillation, where we fine-tune our SWEs (= student) to reproduce the sentence similarities calculated by ST (= teacher). The teacher model is frozen during distillation and only SWEs are fine-tuned; therefore, it works with modest computational resources. Besides, this refinement step does not require labelled data, unlike WordLlama.
Formally, for a mini-batch of $K=128$ sentences, we minimise the following loss $\mathcal{L}$:
@@
where $\mathbf{T},\mathbf{S}\in\mathbb{R}^{K\times K}$ denote the similarity matrices calculated by the teacher and student models, with the value at the $i$-th row and $j$-th column corresponding to the cosine similarity between the $i$-th and $j$-th sentences in the mini-batch. The temperature $\tau$ is set to $0.05$ as commonly done in contrastive learning Gao et al. (2021). [^3] We did not extensively tune this hyper-parameter. Note that the student model represents each sentence by averaging SWEs, while the teacher model encodes them with Transformers and applies its default pooling method. Since the teacher and student models calculate $T$ and $S$ independently, their embedding dimensions can be different and that makes it possible to fine-tune the dimensionality-reduced embeddings with PCA. In contrast, Model2Vec minimises the cosine distance of embeddings output by ST and SWEs, requiring both models to share the same embedding space. Consequently, it applies (word-level) PCA after fine-tuning, and then needs to apply the SIF heuristic to achieve optimal performance. In Section 5.3, we will show that our method yields a similar effect to SIF but in a more refined manner.
For cross-lingual (bilingual) SWEs, we find that performing knowledge distillation for each language leads to worse results on cross-lingual tasks, as it does not encourage SWEs to be mapped into the same cross-lingual space. Therefore, we instead make use of a parallel corpus as a source of knowledge, and fine-tune the cross-lingual SWEs in languages $\ell 1$ and $\ell 2$ by minimising the following loss $\mathcal{L}_{CL}$:
@@ 4.1 Our Models
We train our monolingual (English) SWEs using GTE-base Li et al. (2023), which achieves strong performance on Massive Text Embedding Benchmark (MTEB) among BERT-sized models. We sample sentences from a large-scale unlabelled corpus (CC-100 Wenzek et al. (2020); Conneau et al. (2020)) and use them in each step of our method. Since CC-100 contains paragraphs as well as sentences, we segment text into sentences using a sentence splitter [^5] We use the one at https://github.com/microsoft/BlingFire. and retrieve $N$ short sentences for each word $w\in V$. [^6] We find that using long text for extracting and fine-tuning our SWEs results in worse performance, likely due to the gap of information accessible to the teacher and student models during knowledge distillation (i.e. word order information). For the vocabulary $V$, we include the 150k most frequent words (case-insensitive) in CC-100. To generate cross-lingual SWEs, we use mGTE Zhang et al. (2024), the multilingual version of GTE trained on 75 languages. We train bilingual embeddings for three language pairs, namely English–German (en–de), English–Chinese (en–zh), and English–Japanese (en–ja). We sample short sentences from large-scale parallel corpora (CC-Matrix Schwenk et al. (2021)) and use them to extract SWEs in Section 3.1 and calculate the cross-lingual loss in Eqn. (4). For en–ja, we also use JParaCrawl v3.0 Morishita et al. (2020, 2022) to augment data. For $V$ in English and German, we choose the 150k most frequent words (case-sensitive) in the corresponding corpora, and 30k most frequent tokens (which can be a subword or consist of multiple words) in Japanese and Chinese. [^7] This is because mGTE’s tokeniser segments sentences into tokens without performing word segmentation in both languages (which have no whitespace word boundary).
In Section 3.2, we reduce the dimensionality to $d^{\prime}\in\{256,512\}$, [^8] In most cases, 99% of the variance can be explained with $d^{\prime}=512$ when the original dimensionality $d$ is $768$. and in Section 3.3 we fine-tune SWEs using Adam Kingma and Ba (2015) with the learning rate of 0.001 for 30k steps, and apply early stopping based on the validation loss; the train and validation data consist of sentences sampled from those used in Section 3.1. During inference, we tokenise the text $z$ into words (or tokens in Japanese and Chinese) and strip punctuation, and average SWEs to represent $z$. When a word is not in $V$, we tokenise it into subwords using (m)GTE’s tokeniser, and truncate it by removing the last subword until the remaining string is found in VV. [^9] For instance, if tokeniser is out of vocabulary, we split it into subwords (e.g. token #ise #r), and strip the last subword(s). Then, if tokenise (or token) is in $V$, we use its embedding.